patient was entered the physiological solution (0,9 % solution NaCl) every day in drops (Vang = 500 ml). How many the mg of ions Na " and Cl' have been entered for 5 days of treatment?
1
Expert's answer
2020-09-17T07:15:59-0400

Vang=500ml

And the concetration of NaCl =0.9%

So, v(NaCl) = 500ml × 0.9/100 = 4.5ml

• Number of moles of NaCl will be

n= v/Vm = (4.5/1000)/22.4 = 0.0002mol


NaCl ----- Na+ + Cl-

So in 500 ml of solution there are 0.0002mol Na+ and 0.0002 mol Cl-


So for 5 days if will be 5 times more

n(Na+) =n(Cl-) = 0.001mol